# Who Pays for Fairness? Rethinking Recourse under Social Burden

### v1

@@ 1 Introduction
The ubiquity of automated decision-making systems has considerably reshaped the landscape of how individuals access crucial resources and are granted important social and economic opportunities, ranging from securing credit to receiving essential public services [9, 42, 40]. In recent years, growing concerns have emerged regarding the fairness of these systems, as well as their societal impact [20, 26, 44]. Notably, instances have been documented in which algorithmic decision support systems produced discriminatory outcomes in domains such as welfare distribution [25] and criminal justice [1]. Moreover, in high-stakes scenarios, it is imperative that whenever a model yields a negative decision, it also provides the affected individual with clear and actionable recourse [28]. Unfortunately, affected individuals usually lack transparency and effective means to contest undesired decisions (e.g., bail denial) or take meaningful steps to improve their situation [^1] For example, independent audits have shown fairness concerns and limited contestability options in nationwide-deployed systems, such as COMPAS for recidivism prediction in the United States [1] and VioGén for gender-based violence risk assessment in Spain [17].. For this reason, the “right to recourse” is slowly emerging as a legal requirement in several legislations, e.g., in the European GDPR and the Artificial Intelligence (AI) Act [18].
@@
Pursuing a given recourse often demands considerable effort from the affected individual. As noted by Ustun et al. [47], the average effort to achieve recourse can vary significantly across sensitive groups, raising important concerns about fairness in recourse. This disparity may arise when (i) the recourse itself differs depending on group membership, or when (ii) the recourse is formally identical, but the real-world effort required to carry it out varies in practice. For example, consider a method that offers suggestions to individuals denied a loan. It might recommend ”Increase your monthly income by \$500” to someone from a privileged group, while advising ”Maintain stable employment for 24 more months” to someone from an underprivileged background. The first recommendation is often more viable for individuals with access to high-paying jobs or stable employment opportunities. In contrast, maintaining long-term employment can be substantially more challenging for individuals in precarious jobs or members of low-income and otherwise underprivileged communities. Further, consider an algorithm suggesting that all rejected students retake standardized tests. It may impose minimal burden on those who can afford test preparation, tutoring, and multiple attempts, while students coming from low social and economic backgrounds often face barriers, including limited resources and time constraints due to many socio-economic factors [45]. Thus, what appears to be group-aware or uniform recommendations can translate into very different levels of effort, reinforcing structural inequalities.
(a) The separate treatment of fairness for predictive and recourse elements of the pipeline.
(b) Holistic metrics expose large disparities.
Figure 1: Conventional metrics to measure fairness in recourse hide disparities. (a) Even in a population balanced by sensitive group membership and ground‑truth outcomes (e.g., true ability to repay a loan), treating prediction and recourse fairness separately, as in the majority of prior work, can mask disparities: if positive classifications are unevenly distributed, one group bears recourse burdens more frequently even when per‑instance recourse costs are equal; misclassification (e.g., false negatives) further amplifies this effect. (b) Conventional cost‑parity metrics suggest balanced recourse costs across groups $S\in\{0,1\}$ (left), but our metric (Equation 2) reveals substantial disparity once group differences in decision and error rates are taken into account (right). The algorithms and dataset (CAU-LIN) are from Von Kügelgen et al. [48].
@@
In light of these concerns, it is paramount to develop metrics to reliably quantify and monitor the fairness of recourse recommendations, along with practical methods for improving these metrics. Existing research has primarily examined unfairness in recourse by analyzing the costs imposed on individuals receiving negative (i.e.,“reject”) classifications, either at the individual or group level [48, 15, 29, 4, 50]. However, we argue that fairness in recourse depends not only on what type of recourse is provided but also on who is more likely to receive the reject decision in the first place. As an illustration, consider two groups of people $S=0$ and $S=1$ having the same number of individuals per group. If one of the groups (e.g. $S=1$) has higher acceptance rate than the other group, many more individuals from $S=0$ receive a “reject” decision. Widely utilized fairness-in-recourse metric – equal expected cost (of negatively predicted individuals) – will not capture the inherent unfairness in the system as we can balance the cost of recourse per individual across two groups (e.g. $10/100=100/1000=0.1$). However, individuals in group $S=0$ are disproportionately classified as negative ($1000$ individuals c.f. $100$) and therefore, a higher number of instances from $S=0$ bear the cost of recourse.
@@
We emphasize that the above illustrated problem is amplified when we consider the ground-truth labels, e.g., an individual’s latent (and unobserved) ability to repay a loan instead of just classifier’s acceptance outcomes. Individuals may be wrongly classified and burdened with corrective actions, despite the error stemming from the decision support system itself. This misallocation of responsibility is troubling, particularly as error rates often differ systematically across sensitive groups, a well-documented challenge in algorithmic fairness [34]. All the issues mentioned above stem from the fact that current popular fairness-in-recourse paradigms do not correctly capture fairness issues in the whole system, which includes both the classifier’s components and the recourse’s components.
@@
This paper presents a unified framework for assessing fairness in algorithmic recourse across the full decision pipeline, from initial classification through recourse recommendation (see Figure 1(a) for an overview). Our framework yields formal connections between prediction fairness and recourse fairness, showing how disparities in classification performance can propagate to recourse recommendations, and vice versa. Further, by leveraging these insights, we introduce a lightweight training method, MISOB (MInimax SOcial Burden), that empirically improves fairness throughout the pipeline, without requiring access to sensitive attributes, in several real-world datasets. Lastly, we argue for moving beyond simple group-gap metrics and developing richer measures that more fully describe fairness guarantees.
@@
Our contributions. Summarizing, we (i) theoretically charaterize the sources of unfairness in algorithmic recourse, establishing formal connections between fairness guarantees in recourse and classification (Section 3), (ii) provide evidence of the limitations of approaches based on the standard equal cost paradigm (Section 3.1), (iii) introduce a novel framework for fair recourse (Algorithm 1) based on the concept of social burden (Section 4), and (iv) assess its performance on real-world datasets (Section 5).
@@
Table 1: Summary of the characteristics of existing works on fairness in recourse. We report the proposed fairness metric, whether it considers the classifier ($f(\mathbf{x})$) or the recourse method ($g_{f}(\mathbf{x})$), and the approach used to enhance fairness; specifically, whether it is agnostic to the classifier, the recourse method, or the sensitive attributes ($\mathcal{S}$). A “-” indicates works that propose only evaluation metrics without introducing any algorithm or methodology improving fairness. Our proposed metrics (Section 3) provide a holistic assessment of all components of the pipeline, whereas our approach (Section 4) improves them by remaining agnostic to the underlying components.
@@
$f(\mathbf{x})$ $g_{f}(\mathbf{x})$ $Y$ $f(\mathbf{x})$ $g_{f}(\mathbf{x})$ $\mathcal{S}$
Gupta et al. [22]
Von Kügelgen et al. [48]
Raimondi et al. [43] - - -
Kuratomi et al. [30] - - -
Kavouras et al. [29] - - -
Bell et al. [4]
@@ 2 Preliminaries and Related Work
Each decision subject (i.e., instance) is represented by the triplet $(\boldsymbol{x},\boldsymbol{s},y)$, where $\boldsymbol{x}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ represents the non-sensitive attributes, $\boldsymbol{s}\in\mathcal{S}\subseteq\mathbb{R}^{m}$ the sensitive demographic attributes (e.g., race, gender, age, marital status) and $y\in\mathcal{Y}$ the class label. For simplicity, we assume a binary classification task with $\mathcal{Y}=\{0,1\}$ (e.g., whether the decision subject is capable of repaying the loan). However, the proposed framework is general and can be extended to multiclass settings. We use $P$ to denote the probability of specific events, and $p$ to refer to probability distributions. The sensitive attribute vector $\boldsymbol{s}$ may be multidimensional, allowing individuals to be characterized by multiple protected attributes simultaneously. This enables the analysis of intersectional groups, e.g., “indigenous women under 30”, which capture more nuanced and realistic socio-economic contexts, often subject to distinct and compounded forms of disadvantage.
Further, let $f:\mathcal{X}\rightarrow\mathcal{Y}$ be a binary classifier whose goal is to assign a label in a set $\mathcal{Y}$ to instances in a set $\mathcal{X}$. Let $\delta((\boldsymbol{x},\boldsymbol{s}),\boldsymbol{x}^{\prime})$ denote the cost for an individual with features $(\boldsymbol{x},\boldsymbol{s})$ to change its non-sensitive profile to $\boldsymbol{x}^{\prime}$, e.g., $\ell_{1}$-norm [28]. Given an instance that received a negative classification (i.e., $f(\boldsymbol{x})=0$), we want to find the “closest” instance $\mathbf{x}^{\prime}$ achieving a positive classification by solving the following optimization problem:
$\mathop{\mathrm{argmin}}_{\mathbf{x^{\prime}}\in\mathcal{X}}\;\delta((\boldsymbol{x},\boldsymbol{s}),\mathbf{x}^{\prime})\quad\text{s.t.}\;f(\mathbf{x^{\prime}})\neq f(\mathbf{x})$ (1)
In the literature, $\mathbf{x}^{\prime}$ is also called counterfactual explanation [49]. In our paper, we assume the recommendation $\mathbf{x}^{\prime}=g_{f}(\boldsymbol{x})$ is generated by a recourse algorithm $g_{f}:\mathcal{X}\rightarrow\mathcal{X}$ tailored to the classifier $f$. In algorithmic recourse, a growing body of work has proposed diverse strategies $g_{f}(\boldsymbol{x})$ to offer recourse to individuals assigned unfavorable predictions. These methods vary in several key aspects, including the type of base classifier [47], the optimization technique [11, 36], and the incorporation of actionability or plausibility constraints [27, 10]. Our discussion here is clearly limited; thus, please refer to Karimi et al. [28] for a comprehensive survey on the topic. While providing recommendations is an important step toward accountability and transparency, significant fairness challenges remain, some stemming directly from the recourse process. For example, recourse can exacerbate social segregation [19], and its cost may vary with user preferences [10, 16] and sensitive attributes [50], potentially deepening existing disparities.
@@
Research on algorithmic fairness has primarily examined disparities in model predictions across population subgroups, focusing on the fairness of predictions themselves. Fairness definitions typically fall into two categories: individual-level, which requires similar individuals to receive similar outcomes, and group-level, which evaluates whether performance metrics (e.g., accuracy, true positive rate (TPR)) are comparable across demographic groups. Group metrics are widely used due to their computational simplicity, ease of integration into machine-learning pipelines, and ability to reveal structural disparities. Two common examples are statistical parity [14], which measures differences in the probability of a positive prediction (i.e., acceptance rate (AR)) across groups regardless of true labels, and equality of opportunity [24], which measures such differences only among individuals in the positive class (i.e., TPR). However, these metrics focus solely on the prediction stage and overlook the downstream consequences of model decisions.
@@
Fairness concerns in algorithmic recourse prompted the search for specific fairness criteria [47, 22]. Table 1 summarizes the characteristics of the most relevant approaches that address fairness issues in recourse. Many works proposed both group- and individual-level fairness metrics in this direction [48, 29, 4] and extended the analysis to incorporate classifier predictions and standard predictive performance measures [30, 43]. Several papers that introduced fairness metrics for evaluating algorithmic recourse have also proposed methods to improve fairness according to these metrics [22, 48, 4]. For instance, Von Kügelgen et al. [48] proposed a causality-based method that requires access to a structural causal model (SCM). Gupta et al. [22] introduced an approach applicable only to recourse methods that map instances directly to the decision boundary, while Bell et al. [4] addressed fairness under resource constraints, a setting closer to a ranking rather than a classification problem.
Most fairness metrics proposed in this context focus exclusively on the recourse stage, considering only negatively classified instances and ignoring the classifier’s overall behavior, such as group-specific acceptance and rejection rates. Furthermore, they generally lack practical strategies for improving fairness. None of these studies adopt an intersectional perspective, and all assume access to sensitive attributes for all individuals at training time, an assumption rarely satisfied in real-world applications. Our work is conceptually aligned with Kuratomi et al. [30] and Raimondi et al. [43], but their formulation lacks a clear connection to existing fairness metrics in prediction. Further, these studies neither propose practical algorithms to enhance it nor address the importance of intersectional groups in evaluating recourse fairness.
@@ 3 Revisiting (Un)Fairness in Recourse
In this section, we outline three key limitations of common fairness metrics in pipelines with recourse: (1) they ignore the classifier’s decision behavior, (2) they omit ground truth information, and (3) they focus solely on metric equalization. We propose an alternative formulation that addresses these issues, offers a broader view of unfairness sources in recourse, and establishes explicit links between fairness in prediction and recourse. Proofs are provided in Appendix A.
@@ 3.1 The Impact of the Classifier’s Decisions
Let us exemplify the first issue with a simple example of a binary classification task. Consider a population $X$ with a binary target variable $Y\in\{0,1\}$ and two equally represented sensitive groups $S\in\{0,1\}$, such that $P(S=0)=P(S=1)$ and $P(Y=1\mid S=0)=P(Y=1\mid S=1)$. Let $f(x)$ be a classifier approximating the posterior $P(Y\mid X)$, and assume that under $f$ one group has a higher AR, i.e., $P(f(\boldsymbol{x})=1\mid S=1)>P(f(\boldsymbol{x})=1\mid S=0)$. After recourse is provided to negatively classified instances, let us assume that the expected cost of the required changes is equal across groups: $\mathbb{E}[\delta((\mathbf{x},\mathbf{s}),g_{f}(\mathbf{x}))\mid S=1]=\mathbb{E}[\delta((\mathbf{x},\mathbf{s}),g_{f}(\mathbf{x}))\mid S=0]$. According to standard fairness metrics, this situation would be considered fair. But is it truly fair? While the expected recourse cost is equal once offered, individuals in the group with the lower AR are more likely to require recourse in the first place. We argue that this disparity reflects an unfair situation, which existing popular metrics fail to capture. We now introduce an alternative fairness metric that evaluates recourse costs across the entire population, rather than only among negatively classified instances.
@@
Let $f$ be a classifier and $g_{f}$ the recourse algorithm. Let $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))$ be the cost associated to $\boldsymbol{x}$ of applying the recourse provided by $g_{f}$. Further, consider the expected recourse cost of the sensitive group $\boldsymbol{s}$ as $\mathbb{E}_{P(\boldsymbol{X}=\boldsymbol{x}|S=\boldsymbol{s})}[\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))]$. Then, the expected cost in terms of the classifier’s AR i.e., $P(f(\boldsymbol{x})=1\mid S=\mathbf{s})$, can be defined as:
$\mathcal{C}_{f,g}^{\boldsymbol{s}}=\mathbb{E}[\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))]\cdot(1-P(f(\boldsymbol{x})=1\mid S=\mathbf{s}))$ (2)
where the expectation is over $p(X\mid S=\boldsymbol{s},f(\boldsymbol{x})=0)$.
Namely, Equation 2 arises since instances receiving a positive prediction, $P(f(\boldsymbol{x})=1\mid S=\mathbf{s})$, do not modify their non-sensitive features, implying no changes in recourse cost, $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))=0$.
Figure 1(b) empirically proves the shortcomings of traditional fairness recourse metrics, that capture only the left-hand factor of Equation 2, potentially giving a misleading impression of fairness. Given the previous synthetic binary classification task, we computed the original metric from Von Kügelgen et al. [48] (left), averaging cost only over negatively predicted instances, and our revised formulation from Equation 2, which also accounts for the AR (right) of the base classifiers. When accounting for AR across sensitive groups, disparities in expected recourse costs become substantially larger. For instance, with svm, group S=1S=-1 has the lowest cost under the original metric but the highest under the revised one, as a significantly larger proportion of this group receives negative predictions. Therefore, to avoid masking underlying unfairness, we argue that the cost metric must account for acceptance rates to accurately capture true recourse costs across groups.

### v3

# Ainhize Barrainkua1, Giovanni De Toni2, Jose A. Lozano1,3, Novi Quadrianto1,4

@@ Abstract
Codehttps://github.com/abarrainkua/MISOB
Extended versionhttps://arxiv.org/abs/2509.04128
(a) The separate treatment of fairness for predictive and recourse elements.
(b) Holistic metrics expose large disparities.
Figure 1: Conventional metrics to measure fairness in recourse hide disparities. (a) Given a population balanced by sensitive group membership and ground‑truth outcomes (e.g., true ability to repay a loan), treating prediction and recourse fairness separately, as in the majority of prior work, can mask disparities: if positive classifications are unevenly distributed, one group bears recourse burdens more frequently even when per‑instance recourse costs are equal; misclassification (e.g., false negatives) further amplifies this effect. (b) Conventional cost‑parity metrics suggest balanced recourse costs across groups $S\in\{0,1\}$ (left). Our metric (Eqn. 2) reveals substantial disparity once group differences in decision and error rates are taken into account, even with “fair” classifiers, FairSVM (gupta2019equalizing) and CAU-FairSVM (von2022fairness) (right).
@@ 1 Introduction
The ubiquity of automated decision-making systems has considerably reshaped the landscape of how individuals access crucial resources and are granted important social and economic opportunities, ranging from securing credit to receiving essential public services (chouldechova2018case; purificato2023use; perdomo2025difficult). In recent years, growing concerns have emerged regarding the fairness of these systems, as well as their societal impact (grgichlaca2018human; jobin2019global; raji2022fallacy). Notably, instances have been documented in which algorithmic decision support systems produced discriminatory outcomes in domains such as welfare distribution (heikkila2021dutchscandal) and criminal justice (angwin2022machine). Moreover, in high-stakes scenarios, it is imperative that, whenever a model yields a negative decision, it also provides the affected individual with clear and actionable recourse (karimi2022survey). Unfortunately, affected individuals usually lack transparency and effective means to contest undesired decisions (e.g., bail denial) or take meaningful steps to improve their situation [^2] For example, independent audits have shown fairness concerns and limited contestability options in nationwide-deployed systems, such as COMPAS for recidivism prediction in the United States (angwin2022machine) and VioGén for gender-based violence risk assessment in Spain (eticas2022viogen).. For this reason, the “right to recourse” is slowly emerging as a legal requirement in several legislations, e.g., in the European GDPR and the Artificial Intelligence (AI) Act (act2021proposal).
@@
Pursuing a given recourse often demands considerable effort from the affected individual. As noted by ustun2019actionable, the average effort to achieve recourse can vary significantly across sensitive groups, raising important concerns about fairness in recourse. This disparity may arise when (i) the recourse itself differs depending on group membership, or when (ii) the recourse is formally identical, but the real-world effort required to carry it out varies in practice. For example, consider a method that offers suggestions to individuals denied a loan. It might recommend ”Increase your monthly income by \$500” to someone from a privileged group, while advising ”Maintain stable employment for 24 more months” to someone from an underprivileged background. The first recommendation is often more viable for individuals with access to high-paying jobs or stable employment opportunities. In contrast, maintaining long-term employment can be substantially more challenging for individuals in precarious jobs or members of low-income and otherwise underprivileged communities. Further, consider an algorithm suggesting that all rejected students retake standardized tests. It may impose minimal burden on those who can afford test preparation, tutoring, and multiple attempts, while students coming from low social and economic backgrounds often face barriers, including limited resources and time constraints, due to many socio-economic factors (rodriguez2020socio). Thus, what appears to be group-aware or uniform recommendations can translate into very different levels of effort, reinforcing structural inequalities.
@@
In light of these concerns, it is paramount to develop metrics to reliably quantify and monitor the fairness of recourse recommendations, along with practical methods for improving these metrics. Existing research has primarily examined unfairness in recourse by analyzing the costs imposed on individuals receiving negative classifications (i.e.,“reject”), either at the individual or group level (von2022fairness; ehyaei2023robustness; kavouras2023fairness; bell2024fairness; yetukuri2024providing). However, we argue that fairness in recourse depends not only on what type of recourse is provided but also on who is more likely to receive the negative decision in the first place. As an illustration, consider two groups of people, $S=0$ and $S=1$, having the same number of individuals per group. If one of the groups (e.g., $S=1$) has a higher acceptance rate than the other group, many more individuals from $S=0$ receive a “reject” decision. Widely utilized fairness-in-recourse metric – equal expected cost (of negatively predicted individuals) – will not capture the inherent unfairness in the system as we can balance the cost of recourse per individual across two groups (e.g., $10/100=100/1000=0.1$). However, individuals in group $S=0$ are disproportionately classified as negative ($1000$ individuals vs. $100$) and therefore, a higher number of instances from $S=0$ bear the cost of recourse.
@@
We emphasize that the above illustrated problem is amplified when we consider the ground-truth labels, e.g., an individual’s latent (and unobserved) ability to repay a loan instead of just the classifier’s acceptance outcomes. Individuals may be wrongly classified and burdened with corrective actions, despite the error stemming from the decision support system itself. This misallocation of responsibility is troubling, particularly as error rates often differ systematically across sensitive groups, a well-documented challenge in algorithmic fairness (mehrabi2022surveyfairness). All the issues mentioned above stem from the fact that current popular fairness-in-recourse paradigms do not correctly capture fairness issues in the whole system, which includes both the classifier’s components and the recourse’s components.
@@
This paper presents a unified framework for assessing fairness in algorithmic recourse across the full decision pipeline, from initial classification through recourse recommendation (see Fig. 1(a) for an overview). Our framework yields formal connections between prediction fairness and recourse fairness, showing how disparities in classification performance can propagate to recourse recommendations, and vice versa. Further, by leveraging these insights, we introduce a lightweight training method, MISOB (MInimax SOcial Burden), that empirically improves fairness throughout the pipeline, without requiring access to sensitive attributes, in real-world datasets. Lastly, we argue for moving beyond simple group-gap metrics and developing richer measures that more fully describe fairness guarantees.
Our contributions.
@@
Summarizing, we (i) theoretically characterize the sources of unfairness in algorithmic recourse, establishing formal connections between fairness guarantees in recourse and classification (Section 3), (ii) provide evidence of the limitations of approaches based on the standard equal cost paradigm (Section 3.1), (iii) introduce a novel framework for fair recourse (Algorithm 1) based on the concept of social burden (Section 4), and (iv) assess its performance on real-world datasets (Section 5).
@@
Table 1: Summary of the existing works on fairness in recourse. We report the proposed fairness metric, whether it considers the classifier ($f(x)$) the recourse method ($g_{f}(x)$) or the ground truth outcomes ($Y$), and the approach used to enhance fairness; specifically, whether it is agnostic to the classifier, the recourse method, or the sensitive attributes ($\mathcal{S}$). A “-” indicates works that propose only evaluation metrics without introducing any algorithm or methodology improving fairness.
@@
$f(x)$ $g_{f}(x)$ $Y$ $f(x)$ $g_{f}(x)$ $\mathcal{S}$
gupta2019equalizing
von2022fairness
raimondi2022equality - - -
kuratomi2022measuring - - -
kavouras2023fairness - - -
bell2024fairness
@@ 2 Preliminaries and Related Work
Each decision subject (i.e., instance) is represented by the triplet [^3] We denote random variables with capital letters, their realizations with lowercase letters, and sets with capital letters in italic. $(x,s,y)$, where $x\in\mathcal{X}\subseteq\mathbb{R}^{d}$ represents the non-sensitive attributes, $s\in\mathcal{S}\subseteq\mathbb{R}^{m}$ the sensitive demographic attributes (e.g., race, gender, age), and $y\in\mathcal{Y}$ the class label. For simplicity, we assume a binary classification task with $\mathcal{Y}=\{0,1\}$ (e.g., whether the decision subject is capable of repaying the loan). However, the proposed framework is general and can be extended to multiclass settings. The sensitive attributes $s$ can be multidimensional, enabling the characterization of individuals by multiple protected attributes and the analysis of intersectional groups (e.g., indigenous women under 30) that reflect more nuanced and compounded socio-economic disadvantages. Further, let $f:\mathcal{X}\rightarrow\mathcal{Y}$ be a binary classifier whose goal is to assign a label in a set $\mathcal{Y}$ to instances in a set $\mathcal{X}$. Let $\delta((x,s),x^{\prime})$ denote the cost for an individual with features $(x,s)$ to change its non-sensitive profile to $x^{\prime}$, e.g., $\ell_{1}$-norm (karimi2022survey). Given an instance that received a negative classification (i.e., $f(x)=0$), we want to find the “closest” instance $x^{\prime}$ achieving a positive classification by solving the following optimization problem:
$\mathop{\mathrm{argmin}}_{x^{\prime}\in\mathcal{X}}\;\delta((x,s),x^{\prime})\quad\text{s.t.}\;f(x^{\prime})\neq f(x)$ (1)
In the literature, $x^{\prime}$ is also called counterfactual explanation (wachter2017counterfactual). Here, we assume the recommendation $x^{\prime}=g_{f}(x)$ is generated by a recourse algorithm $g_{f}:\mathcal{X}\rightarrow\mathcal{X}$ tailored to the classifier $f$. A growing body of work has proposed diverse strategies $g_{f}(x)$ to offer recourse to individuals assigned unfavorable predictions. These methods vary in several key aspects, including the type of base classifier considered, and actionability constraints. Our discussion here is clearly limited; thus, please refer to karimi2022survey for a comprehensive survey on the topic. While providing recommendations is an important step toward accountability and transparency, significant fairness challenges remain, some stemming directly from the recourse process. Indeed, recourse can exacerbate social segregation (gao2023impact), and its cost may vary with user preferences (detoni2023synrec; detoni2024personalized) and sensitive attributes (yetukuri2024providing), potentially deepening existing disparities.
@@
Research on algorithmic fairness has primarily examined disparities in model predictions across population subgroups, focusing on the fairness of predictions themselves. Fairness definitions typically fall into two categories: individual-level, which requires similar individuals to receive similar outcomes, and group-level, which evaluates whether performance metrics (e.g., accuracy) are comparable across demographic groups. Group metrics are widely used due to their computational simplicity, ease of integration into machine-learning pipelines, and ability to reveal structural disparities. Two common examples are statistical parity (dwork2012fairness), which measures differences in the probability of a positive prediction (i.e., acceptance rate) across groups regardless of true labels, and equality of opportunity (heidari2019moral), which measures such differences only among individuals in the positive class (i.e., true positive rate). However, these metrics focus solely on the prediction stage and overlook the downstream consequences of model decisions.
@@
Fairness concerns in algorithmic recourse prompted the search for specific fairness criteria (ustun2019actionable; gupta2019equalizing). Table 1 summarizes the characteristics of the most relevant approaches that address fairness issues in recourse. Many works proposed both group- and individual-level fairness metrics in this direction (von2022fairness; kavouras2023fairness; bell2024fairness) and extended the analysis to incorporate classifier predictions and standard predictive performance measures (kuratomi2022measuring; raimondi2022equality). Several papers that introduced fairness metrics for evaluating algorithmic recourse have also proposed methods to improve fairness according to these metrics (gupta2019equalizing; von2022fairness; bell2024fairness). For instance, von2022fairness proposed a causality-based method that requires access to a structural causal model (SCM). gupta2019equalizing introduced an approach applicable only to recourse methods that map instances directly to the decision boundary, while bell2024fairness addressed fairness under resource constraints, a setting closer to a ranking rather than a classification problem.
Most fairness metrics proposed in this context focus exclusively on the recourse stage, considering only negatively classified instances and ignoring the classifier’s overall behavior, such as group-specific acceptance and rejection rates. Furthermore, they generally lack practical strategies for improving fairness. Lastly, none of these studies adopt an intersectional perspective, and all assume access to sensitive attributes for all individuals at training time, an assumption rarely satisfied in real-world applications. Our work is conceptually aligned with kuratomi2022measuring and raimondi2022equality, but their formulation lacks a clear connection to existing fairness metrics in prediction.
@@ 3 Revisiting (Un)Fairness in Recourse
We now outline three key limitations of common fairness metrics in pipelines with recourse: (1) they ignore the classifier’s decision behavior, (2) they omit ground truth information, and (3) they focus solely on metric equalization. We propose an alternative formulation that addresses these issues, offers a broader view of unfairness sources in recourse, and establishes explicit links between fairness in prediction and recourse. Proofs are provided in Appendix A.
@@ 3.1 The Impact of the Classifier’s Decisions
Let us exemplify the first issue with an example of a binary classification task. Consider a population $X$ with a binary target variable $Y\in\{0,1\}$ and two equally represented sensitive groups $S\in\{0,1\}$, such that $P(S=0)=P(S=1)$ and $P(Y=1\mid S=0)=P(Y=1\mid S=1)$. Let $f(x)$ be a classifier approximating the posterior $P(Y\mid X)$, and assume that under $f$ one group has a higher acceptance rate (AR), $P(f(x)=1\mid S=1)>P(f(x)=1\mid S=0)$. After recourse is provided to negatively classified instances, let us assume that the expected cost of the required changes is equal across groups $\mathbb{E}[\delta((x,s),g_{f}(x))\mid S=1]=\mathbb{E}[\delta((x,s),g_{f}(x))\mid S=0]$. According to standard fairness metrics, this situation would be considered fair. But is it truly fair? While the expected recourse cost is equal once offered, individuals in the group with lower AR are more likely to require recourse in the first place. We argue that this disparity reflects an unfair situation, which existing popular metrics fail to capture. We now introduce an alternative fairness metric that evaluates recourse costs across the entire population, rather than only among negatively classified instances.
@@
Let $f$ be a classifier and $g_{f}$ a recourse algorithm. Given a sensitive group $s\in\mathcal{S}$, let $\delta((x,S=s),g_{f}(x))$ be the cost of applying the recourse provided by $g_{f}$. Then, the expected recourse cost $C_{f,g}^{s}$ can be defined as:
$\underbrace{\mathbb{E}[\delta((X,S=s),g_{f}(X))]}_{\text{Expected cost for instances with f(x)=0}}(1-\underbrace{P(f(X)=1\mid S=s)}_{\text{Acceptance Rate (AR)}})$ (2)
where the expectation is over $P(X\mid S=s,f(X)=0)$, and $P(f(X)=1\mid S=s)$ indicates the acceptance rate of the classifier for the given sensitive group.
Namely, Eqn. 2 arises since instances receiving a positive prediction, $\{x:f(x)=1,\;x\in X\}$, do not modify their non-sensitive features, implying no changes in recourse cost, $\delta((x,s),g_{f}(x))=0$. Fig. 1(b) shows how, when accounting for AR across sensitive groups, disparities in expected recourse costs become substantially larger. Given the CAU-LIN dataset from von2022fairness, we computed the empirical expected recourse cost by simply averaging only over negatively predicted instances (left), and by using Eqn. 2 (right), which also accounts for the AR of the base classifiers. Indeed, Fig. 1(b) shows that both groups have similar costs under the original metric with FairSVM, but $S=1$ has the highest under the revised one, as a significantly larger proportion of this group receives negative predictions. Therefore, to avoid masking underlying unfairness, we argue that the cost metric must account for acceptance rates to accurately capture true recourse costs across groups.